OpenStudy (anonymous):

$17,889 is invested , part at 8% and the rest at 5% if the interest earned from the amount invested at 8% exceeds the interest earned from the amount invested at 5% by $412.44, how much is invested at each rate? (round to two decimals place if necessary) ....use x and y to set up a system of equation to solve for the given problem

OpenStudy (anonymous):

0.08x-0.05y=41244 x+y=17889 so x= 17880-y so now it is simultaneous you can substitute x from second equation into first and find y then substitute in second equation to find x 0.08(17880-y) -0.05y = 41244 when you found y substitute to find x

OpenStudy (anonymous):

thank you! I got x = 17888 and y =20372 @anette

OpenStudy (anonymous):

great